# ShapeICP: Iterative Category-Level Object Pose and Shape Estimation from Depth

### v2

@@ Abstract
Category-level object pose and shape estimation from a single depth image has recently drawn research attention due to its potential utility for tasks such as robotics manipulation. The task is particularly challenging because the three unknowns, object pose, object shape, and model-to-measurement correspondences, are compounded together, but only a single view of depth measurements is provided. Most of the prior work heavily relies on data-driven approaches to obtain solutions to at least one of the unknowns, and typically two, running with the risk of failing to generalize to unseen domains. The shape representations used in the prior work also mainly focus on point cloud and signed distance field (SDF). In stark contrast to the prior work, we approach the problem using an iterative estimation method that does not require learning from pose-annotated data. In addition, we adopt a novel mesh-based object active shape model that the previous literature has not explored. Our algorithm, ShapeICP, is based on the iterative closest point (ICP) algorithm but is equipped with additional features for the category-level pose and shape estimation task. Although not using pose-annotated data, ShapeICP surpasses many data-driven approaches that rely on pose data for training, opening up a new solution space for researchers to consider.
@@ I Introduction
The category-level object pose and shape estimation problem is typically set up such that the instance segmentation and semantic classification of the object are performed on the accompanying RGB image with an off-the-shelf method [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. However, even after factoring out the object segmentation problem, inferring the object pose and shape from the segmented depth image remains challenging due to the three entangled unknowns: object pose, object shape, and model-to-measurement correspondences. It requires us to know two of the three unknowns to reduce the problem to an easily solvable one. For instance, if we know the object shape and how the measurements correspond to the points on the shape, it is reduced to the problem of aligning two point sets, which the Umeyama algorithm [13] can solve. If the object pose and the correspondences of measurements to a shape template are known, one can transform and deform the template to the measurements to obtain the shape. Researchers often recognize the correspondences also as unknowns and deform the template [14] using Chamfer loss, where the closest points are the correspondences [15, 16, 14]. However, with the three variables all unknown, the problem is extremely difficult.
@@
Most prior methods exploit neural networks to learn a mapping from inputs such as the RGB-D image and a categorical shape prior to one or two of the three unknowns. Although these learning methods greatly factor away the complexities in the problem by leveraging data and have achieved excellent performance on benchmarks, they face potential failures caused by domain changes and the tedious data curation process required for training when they are deployed to a new environment. The ground truth pose of an object is especially painstaking to annotate since so far there has been no scalable and easily accessible approach to measure the pose of an object. On the other hand, virtually all the prior methods focus on either point cloud or signed distance field (SDF) as the shape representation. Point clouds are flexible but do not have the notion of surface, while SDF entails unfavorable computational complexity, which is often scaled cubically with the resolution.
In this work, we take a radically different approach. First, in contrast to prior methods, we adopt a novel mesh-based active shape model (ASM) [17] as the object shape representation. Meshes incorporate the notion of surface, being superior to point clouds. They are also inherently a geometric representation, unlike an SDF or occupancy grid where the geometry is represented implicitly and some processing, such as finding the zero level set (SDF) or marching cubes [18] (occupancy grid), is required to extract the geometry. Using meshes in category-level object pose and shape estimation is also a relatively untapped terrain in the literature; it is worth exploring more. Second, although methods such as self-supervised learning and sim-to-real transfer can help circumvent the annotation difficulty, we seek a different avenue by devising an iterative estimation algorithm based on the iterative closest point (ICP) algorithm [19] and many other classical estimation techniques without training any module on the pose-annotated data. The method has the option of using an image-based shape classification network for shape initialization to boost estimation performance. We consider that in practice, the ground-truth shape of an object is more accessible than its ground-truth pose since object scanning and depth fusion techniques [20] are mature.
@@
We evaluate our method on the NOCS REAL dataset [21] and find it better or comparable to many learning-based methods, even though our method is at the disadvantage that no pose-annotated data are used. The evaluation of our method also reveals what can already be accomplished without learning and why learning may be needed in some places. These insights are traditionally difficult to draw from the evaluation of black-box style learning-based methods.
@@ II-A Category-Level Pose and Shape Estimation
Predicting shape and pose. CASS [4] first learns a point cloud based shape embedding space. During inference, the shape latent code is predicted given the RGB-D image patch. The predicted latent code is both decoded to the shape and fed to the module for pose prediction. CenterSnap [22] and ShAPO [23] build a YOLO-style [24] one-shot estimation pipeline. iCaps [5] selects the rotation by comparing the depth image features to the features in a pre-computed code book and predicts the DeepSDF [25] latent vector as the shape representation. Other methods that directly regress shape and pose to various extents include [6, 7, 26].
@@
Predicting pose. This category of methods bypasses shape estimation and estimates only the pose. FS-Net [27] extracts rotation-aware features through a scale and shift-invariant 3D graph convolution network. It decouples the rotation into two perpendicular vectors for prediction. SAR-Net [8] trains neural networks to transform a categorical template to the same rotation as the observed point cloud and complete the observed point cloud by mirroring about the symmetry plane. These network outputs are used to compute the pose.
@@
Predicting correspondences. [28] assumes known correspondences provided by neural keypoint matching and solves an alignment problem between the model keypoints and the measurement keypoints. Instead of only estimating the $\mathrm{SIM}(3)$ pose as in Umeyama [13], the method additionally estimates the parameters for an active shape model (ASM) [17] of the object. The ASM is linear and can be treated similarly to translation. The method also features a graph-theoretic approach and an outlier-robust solver to handle false correspondences. [29] is an earlier work that essentially shares the same problem formulation as [28] but with a different solver that does not consider the outlier issue.
@@
Optimization-based. Neural Analysis-by-Synthesis [9] leverages neural rendering to render an image from the object shape code and pose. The rendered image is compared with the actual image to optimize the shape code and pose. DirectShape [10] sets up the optimization losses by silhouette matching and left-right stereo photo consistency using an SDF-based ASM. Strong regularization, such as the ground plane constraint, is engaged to alleviate the optimization difficulty. ELLIPSDF [12] employs a bi-level shape representation to coarsely fit the object segmentation for initialization and then finely optimize the pose and shape.
@@
As we have seen, previous methods commonly adopt data-driven solutions. Pose-annotated data are exploited in almost all the prior work, either explicitly or implicitly. Even the optimization-based methods may actually require extra information such as multiple views [12] or application-specific constraints [10]. Our method works completely without pose-annotated data. The addition of an optional shape classification network merely requires shape-annotated data, which also falls into the rare category of predicting only the shape, a direction not well explored by the prior work.
@@ II-B Shape Representation – Active Shape Model
The active shape model (ASM) [17] has been employed in the category-level object pose and shape estimation task [10, 29, 28, 6], and also in the 3D object shape reconstruction task [30]. The ASM has the advantage of being a simple linear model that is more efficient to back-propagate through during optimization compared to a neural network. The object ASM is primarily point-based [29, 28, 6] and SDF-based [10, 30]. It is interesting to explore the mesh representation further.
@@ III-A Overview
Our task is to estimate the $\mathrm{SIM}(3)$ pose and shape of an object which is only known up to its semantic category given a single depth image of it. We assume that instance segmentation of the object and its class label are provided by an off-the-shelf method such as Mask RCNN [31] as typically used in the prior work [9, 10, 1, 2, 3, 4, 6, 5, 7, 8]. We adopt a mesh-based object active shape model (ASM) [17] for its flexibility and surface representability. After obtaining the ASM offline (Section III-B), we estimate the pose and shape by transforming and fitting the shape model to the observed depth points. The optimization is realized by an augmented iterative closest point (ICP) algorithm which, in addition to the existing pose estimation step, has a shape deformation step for shape estimation (Section III-C). We thus name our algorithm ShapeICP. However, a significant challenge is that ICP is a local solver, and so is the base form of ShapeICP. To overcome this, ShapeICP uses multi-hypothesis tracking for rotation estimation, expectation maximization (EM) for robust correspondence handling, and shape classification for shape initialization (Section III-D). The full pseudo-algorithm of ShapeICP can be found in Section III-E.
@@ III-B Mesh-based Object Active Shape Model
Template deformation. Mesh-based ASMs have been applied to human faces [32] but not yet in the task of object pose and shape estimation. The bottleneck is that object models from a database have different numbers of vertices and different vertex connectivity even within the same category, making it difficult to draw corresponding vertices across models for PCA. To solve this challenge, we deform a template mesh to wrap around each object model and run PCA on the vertices of the deformed templates for all the object models within a category. Specifically, we denote a mesh by $\mathcal{M=(V,E,F)}$, where $\mathcal{V}=\{v_{i}\}_{i=1}^{V}$ are the $V$ vertices in the mesh, =$\mathcal{E}=\{e_{i}\}_{i=1}^{E}$ is the set of $E$ tuples of the indices of a pair of vertices connected by an edge, =$\mathcal{F}=\{f_{i}\}_{i=1}^{F}$ is the set of $F$ tuples of indices of vertices on the same face. To deform the template (source) mesh $\mathcal{M}_{s}$ to a target mesh $\mathcal{M}_{t}$, we randomly sample points on the mesh surface with $N$ points from the template mesh $p_{n}\sim\mathcal{M}_{s}$ and $M$ points from the target mesh $q_{m}\sim\mathcal{M}_{t}$ (here $N=M$). The deformation is done by minimizing a similar set of losses as in [16, 15, 33]:
@@
where $\mathcal{N}(i)=\{j\}~{}\forall(i,j)\in\mathcal{E}$ is the set of neighboring vertices of vertex $v_{i}$. The overall objective function is $\min_{\mathcal{V}_{s}}\mathcal{L}_{c}+\lambda_{n}\mathcal{L}_{n}+\lambda_{e}\mathcal{L}_{e}+\lambda_{l}\mathcal{L}_{l}$, where $\lambda$’s are the weights for the loss terms. This optimization is carried out by stochastic gradient descent (SGD) [34] starting from a spherical template with $p_{n}$ and $q_{m}$ resampled in each step. An example of the process is visualized in Fig. 1.
Figure 1: An example of the template deformation process with the initial template on the left and the target mesh on the right side of the arrow.
PCA. Now that we have acquired a set of deformed templates that take on the model shapes across a category and have a consistent number of vertices and connectivity, we can build a feature vector for each model by concatenating the vertices of its deformed template, i.e. $[v_{1}...v_{V_{s}}]$, and then running PCA on all the feature vectors to obtain our category-level mesh-based ASM. A vertex in the final ASM with $K$ principal components is expressed as
@@
where $b_{0,i}$ is the vertex in the mean shape corresponding to $v_{i}$, $b_{k,i}$ is the corresponding vertex in basis $k$, and $\{c_{k}\}_{k=1}^{K}$ are the weights for the bases. These weights parameterize the shape. Fig. 2 visualizes the vertices of the bases and the final shape in an example ASM. The edges and faces of the mesh-based ASM are inherited from the template mesh.
@@ III-C Alternating Pose and Shape Optimization
where $\mathsf{F}(n)$ is a function that maps the sample index $n$ to a face randomly selected with probability proportional to the face area, and $\mu_{n,i}$ are randomly generated interpolation weights for this sample. These interpolation weights are fixed during the later optimization. Combining with (5), $p_{n}=\sum_{i\in\mathsf{F}(n)}{\mu_{n,i}(b_{0,i}+\sum_{k=1}^{K}{c_{k}b_{k,i}})}$. We now introduce the ShapeICP formulation:
@@
where $\{q_{m}\}_{m=1}^{M}$ (overloaded notation) are now the object depth points segmented from the depth image and back-projected to 3D using the camera intrinsics, $R$, $t$, and $s$ are the object rotation, translation, and scale to be estimated, $\mathbf{c}$ which enters the objective function through $p_{n}$ is the ASM shape code also to be estimated. For each measurement $q_{m}$, (7) encourages the transformed model points to be close to the measurements, which is similar to the ICP objective function [19] but with the addition of the shape parameter $\mathbf{c}$. We also remark that formulation (7) equivalently works for a point-based ASM where $p_{n}$ is simply a point in the ASM point cloud. However, a mesh-based ASM has the notion of surface and thus is a better geometric representation. Furthermore, a mesh model enables the possibility of rendering to be used in Section III-D.
Core algorithm. We make two observations of (7). First, if the shape code $\mathbf{c}$ is given, the problem shares the same objective function as ICP. Second, if the $\mathrm{SIM}(3)$ pose is given, the problem mimics the mesh deformation problem introduced in Section III-B except that (7) is only one-sided Chamfer distance and the optimization variable is the shape code instead of the vertices. These two observations lead to an intuitive alternating minimization algorithm to solve (7):
@@
After the pose step, the incremental $R$, $t$, and $s$ are accumulated onto the estimates $\hat{R}$, $\hat{t}$, and $\hat{s}$. Every time $\mathbf{c}$ is updated, $p_{n}$’s are re-sampled to ensure sample uniformity as the areas of the mesh faces change.
Shape step. In the shape step, we include the regularization losses (2) – (4). The final objective function $\mathcal{L}_{sc}$ ($sc$: shape code) in the shape step is $\min_{\mathbf{c}}{\mathcal{L}_{sc}}=\min_{\mathbf{c}}\mathcal{L}_{ps}+\lambda_{n}\mathcal{L}_{n}+\lambda_{e}\mathcal{L}_{e}+\lambda_{l}\mathcal{L}_{l}$, where $\mathcal{L}_{ps}$ ($ps$: pose shape) refers to the objective function in (7) and $\lambda$’s are the weights. These weights can be reused from Section III-B if we center and normalize the observed point cloud, and shift and normalize the transformed model points by the same amount (because the ShapeNet [35] objects in Section III-B are all centered and normalized). Since this is an iterative coordinate-descent style algorithm, we prefer small increments on $\mathbf{c}$. Only one step of gradient descent is performed on $\mathbf{c}$ during each shape step. However, we repeat the shape step several times in each iteration. The shape step is skipped in the final iterations for better empirical performance.

### v3

@@ Abstract
Category-level object pose and shape estimation from a single depth image has recently drawn research attention due to its potential utility for tasks such as robotics manipulation. The task is particularly challenging because the three unknowns, object pose, object shape, and model-to-measurement correspondences, are compounded together, but only a single view of depth measurements is provided. Most of the prior work heavily relies on data-driven approaches to obtain solutions to at least one of the unknowns, and typically two, risking generalization failures if not designed and trained carefully. The shape representations used in the prior work also mainly focus on point clouds and signed distance fields (SDFs). In stark contrast to the prior work, we approach the problem using an iterative estimation method that does not require learning from pose-annotated data. Moreover, we construct and adopt a novel mesh-based object active shape model (ASM), which additionally maintains vertex connectivity compared to the commonly used point-based object ASM. Our algorithm, ShapeICP, is based on the iterative closest point (ICP) algorithm but is equipped with additional features for the category-level pose and shape estimation task. Although not using pose-annotated data, ShapeICP surpasses many data-driven approaches that rely on pose data for training, opening up a new solution space for researchers to consider.
@@ I Introduction
The category-level object pose and shape estimation problem is typically set up in a way that the instance segmentation and semantic classification of the object are performed on the accompanying RGB image with an off-the-shelf method [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. However, even after factoring out the object segmentation problem, inferring the object pose and shape from the segmented depth image remains challenging due to the three entangled unknowns: object pose, object shape, and model-to-measurement correspondences. It requires us to know two of the three unknowns to reduce the problem to an easily solvable one. For instance, if we know the object shape and how the measurements correspond to the points on the shape, it is reduced to the problem of aligning two point sets, which the Umeyama algorithm [13] can solve. If the object pose and the correspondences between the measurements and a shape template are known, one can transform and deform the template to fit the measurements to obtain the shape. Researchers often recognize the correspondences also as unknowns and deform the template [14] using the Chamfer loss, where the closest points are the correspondences [15, 16, 14]. However, with the three variables all unknown, the problem is extremely difficult.
@@
Most prior methods exploit neural networks to learn a mapping from inputs such as the RGB-D image and a categorical shape prior to one or two of the three unknowns. Although these learning methods greatly factor away the complexities in the problem by leveraging data and have achieved excellent performance on benchmarks, they face potential failures caused by domain changes and the tedious data curation process required for training when they are deployed to a new environment. The ground truth pose of an object is especially painstaking to annotate since so far there has been no scalable and easily accessible approach to measure the pose of an object in the real world. On the other hand, virtually all the prior methods focus on either point clouds or signed distance fields (SDFs) as the shape representation. Point clouds are flexible but lack the notion of surface, while SDFs entail unfavorable computational complexity, which often scales cubically with resolution.
In this work, we take a radically different approach and make the following contributions. First, in contrast to prior methods, we adopt a novel mesh-based active shape model (ASM) [17] (instead of the commonly used point-based or SDF-based object ASM) as the shape representation. Meshes are an explicit geometric representation, unlike SDFs or occupancy grids where the geometry is represented implicitly and some processing, such as finding the zero level set (SDF) or marching cubes [18] (occupancy grid), is required to extract the geometry. A mesh-based ASM also incorporates the notion of surface (or vertex connectivity), being superior to point ASM. However, the addition of the connectivity information requires a different ASM construction methodology, which is also our contribution.
Second, although methods such as self-supervised learning and sim-to-real transfer can help circumvent the annotation difficulty, we seek a different avenue by devising a learning-free iterative estimation algorithm based on the iterative closest point (ICP) algorithm [19] and alternating minimization. Third, we further adapt a set of robust estimation techniques, such as expectation-maximization, multi-hypothesis tracking, and shape classification, to cope with the local minimum problem in the estimation process. The option of using an image-based shape classification network for shape initialization only requires training on the ground-truth object shape, which is practically more accessible than the ground-truth pose since object scanning and depth fusion techniques [20] are mature.
@@
Lastly, we evaluate our method on the NOCS REAL dataset [21] and find it better or comparable to many learning-based methods, even though our method is at the disadvantage that no pose-annotated data are used. The evaluation of our method also reveals what can already be accomplished without learning and why learning may be needed in some places. These insights are traditionally difficult to draw from the evaluation of black-box style learning-based methods. We provide more extensive evaluation and analysis in the Appendix.
@@ II-A Category-Level Pose and Shape Estimation
Predicting shape and pose. CASS [4] first learns a point cloud based shape embedding space. During inference, the shape latent code is predicted given the RGB-D image patch. The predicted latent code is both decoded to the shape and fed to a module for pose prediction. CenterSnap [22] and ShAPO [23] build a YOLO-style [24] one-shot estimation pipeline. iCaps [5] selects the rotation by comparing the depth image features to the features in a pre-computed code book and predicts the DeepSDF [25] latent vector as the shape representation. Other methods that directly regress shape and pose to various extents include [6, 7, 26].
@@
Predicting pose. This category of methods bypasses shape estimation and estimates only the pose. FS-Net [27] extracts rotation-aware features through a scale- and shift-invariant 3D graph convolution network. It decouples the rotation into two perpendicular vectors for prediction. SAR-Net [8] trains neural networks to transform a categorical template to the same rotation as the observed point cloud and complete the observed point cloud by mirroring about the symmetry plane. These network outputs are used to compute the pose.
@@
Predicting correspondences. [28] assumes known correspondences provided by neural keypoint matching and solves an alignment problem between the model keypoints and the measurement keypoints. Instead of only estimating the $\mathrm{SIM}(3)$ pose as in Umeyama [13], the method additionally estimates the parameters for a point-based active shape model (ASM) [17] of the object. The ASM is linear and can be treated similarly to translation. The method also features a graph-theoretic approach and an outlier-robust solver to handle false correspondences. [29] is an earlier work that essentially shares the same problem formulation as [28] but with a different solver that does not consider the outlier issue.
@@
Optimization-based. Neural Analysis-by-Synthesis [9] leverages neural rendering to render an image from the object shape code and pose. The rendered image is compared with the actual image to optimize the shape code and pose. DirectShape [10] sets up the optimization losses by silhouette matching and left-right stereo photo consistency using an SDF-based ASM. Strong regularization, such as the ground plane constraint, is engaged to alleviate the optimization difficulty. ELLIPSDF [12] employs a bi-level shape representation to coarsely fit the object segmentation for initialization and then finely optimize the pose and shape using multiple views.
@@
Summary. As we have seen, previous methods commonly adopt data-driven solutions. Pose-annotated data are exploited in almost all the prior work, either explicitly or implicitly. Even the optimization-based methods may actually require extra information such as multiple views [12] or application-specific constraints [10]. Our method works completely without pose-annotated data. The addition of an optional shape classification network merely requires shape-annotated data, which also falls into the rare category of predicting only the shape, a direction not well explored by prior work.
@@ II-B Shape Representation – Active Shape Model
The active shape model (ASM) [17] has been employed in the category-level object pose and shape estimation task [10, 29, 28, 6], and also in the 3D object shape reconstruction task [30]. The ASM has the advantage of being a simple linear model that is more efficient to back-propagate through during optimization compared to a neural network. The object ASM is primarily point-based [29, 28, 6] and SDF-based [10, 30]. It is interesting to explore the mesh-based ASM further.
@@ III-A Overview
Our task is to estimate the $\mathrm{SIM}(3)$ pose and the shape of an object which is only known up to its semantic category given a single depth image of it. We assume that the instance segmentation of the object and its class label are provided by an off-the-shelf method such as Mask RCNN [31], as is typically done in the prior work [9, 10, 1, 2, 3, 4, 6, 5, 7, 8]. We adopt a mesh-based object active shape model (ASM) [17] for its flexibility and surface representability. After obtaining the ASM offline (Section III-B), we estimate the pose and shape by transforming and fitting the shape model to the observed depth points. The optimization is realized by an augmented iterative closest point (ICP) algorithm which, in addition to the existing pose estimation step, has a shape deformation step for shape estimation (Section III-C). We thus name our algorithm ShapeICP. However, a significant challenge is that ICP is a local solver, and so is the base form of ShapeICP. To overcome this, ShapeICP uses multi-hypothesis tracking for rotation estimation, expectation maximization (EM) for robust correspondence handling, and shape classification for shape initialization (Section III-D). The full pseudo-algorithm of ShapeICP can be found in Section III-E.
@@ III-B Mesh-based Object Active Shape Model
Template deformation. Mesh-based ASMs have been applied to human faces [32] but not yet in the task of object pose and shape estimation. The bottleneck is that object models from a database have different numbers of vertices and different vertex connectivity even within the same category, making it difficult to draw corresponding vertices [6] and maintain consistent vertex connectivity across models for PCA (elaborated in the Appendix). To solve these challenges, we deform a template mesh to wrap around each object model and run PCA on the vertices of the deformed templates for all the object models within a category. Specifically, we denote a mesh by $\mathcal{M=(V,E,F)}$, where $\mathcal{V}=\{v_{i}\}_{i=1}^{V}$ are the $V$ vertices in the mesh, $\mathcal{E}=\{e_{i}\}_{i=1}^{E}$ is the set of $E$ tuples of the indices of a pair of vertices connected by an edge, $\mathcal{F}=\{f_{i}\}_{i=1}^{F}$ is the set of $F$ tuples of indices of vertices on the same face. $\mathcal{E}$ or $\mathcal{F}$ defines the vertex connectivity. To deform the template (source) mesh $\mathcal{M}_{s}$ to a target mesh $\mathcal{M}_{t}$, we randomly sample points on the mesh surface with $N$ points from the template mesh $p_{n}\sim\mathcal{M}_{s}$ and $M$ points from the target mesh $q_{m}\sim\mathcal{M}_{t}$ (here $N=M$). As in [16, 15, 33], the deformation can be done by minimizing losses:
@@
where $\mathcal{N}(i)=\{j\}~\forall(i,j)\in\mathcal{E}$ is the set of neighboring vertices of vertex $v_{i}$. The overall objective function is $\min_{\mathcal{V}_{s}}\mathcal{L}_{c}+\lambda_{n}\mathcal{L}_{n}+\lambda_{e}\mathcal{L}_{e}+\lambda_{l}\mathcal{L}_{l}$, where the $\lambda$’s are the weights for the loss terms. This optimization is carried out by stochastic gradient descent (SGD) [34] starting from a spherical template with $p_{n}$ and $q_{m}$ resampled in each step. An example of the process is visualized in Fig. 1.
[…]
Figure 1: An example of the template deformation process.
PCA. Now that we have acquired a set of deformed templates that take on the model shapes across a category and have a consistent number of vertices and the same connectivity, we can build a feature vector for each model by concatenating the vertices of its deformed template, i.e. $v_{1}...v_{V_{s}}]$, and then running PCA on all the feature vectors to obtain our category-level mesh-based ASM. A vertex in the final ASM with $K$ principal components is expressed as
@@
where $b_{0,i}$ is the vertex in the mean shape corresponding to $v_{i}$, $b_{k,i}$ is the corresponding vertex in the basis $k$, and $\{c_{k}\}_{k=1}^{K}$ are the weights for the bases. These weights parameterize the shape. Fig. 2 visualizes the vertices of the bases and the final shape in an example ASM. The edges ($\mathcal{E}$) and faces ($\mathcal{F}$) (i.e. connectivity) of the mesh-based ASM are inherited from the template mesh.
@@ III-C Alternating Pose and Shape Optimization
where $\mathsf{F}(n)$ is a function that maps the sample index $n$ to a face randomly selected with probability proportional to the face area, and $\mu_{n,i}$ are randomly generated interpolation weights for this sample. These interpolation weights are fixed during the later optimization. Combining with (5), $p_{n}=\sum_{i\in\mathsf{F}(n)}{\mu_{n,i}(b_{0,i}+\sum_{k}{c_{k}b_{k,i}})}$. We now introduce the ShapeICP formulation:
@@
where $\{q_{m}\}_{m=1}^{M}$ (overloaded notation) are now the object depth points segmented from the depth image and back-projected to 3D using the camera intrinsics, $R$, $t$, and $s$ are the object rotation, translation, and scale to be estimated, $\mathbf{c}$ which enters the objective function through $p_{n}$ is the ASM shape code also to be estimated. For each measurement $q_{m}$, (7) encourages the transformed model points to be close to the measurements, which is similar to the ICP objective function [19] but with the addition of the shape parameter $\mathbf{c}$. We also remark that formulation (7) equivalently works for a point-based ASM [6] where $p_{n}$ is simply a point in the ASM point cloud. However, a mesh-based ASM has the notion of surface and thus is a better geometric representation. Furthermore, a mesh model enables the possibility of rendering to be used in Section III-D.
Core algorithm. We make two observations of (7). First, if the shape code $\mathbf{c}$ is given, the problem shares the same objective function as ICP. Second, if the $\mathrm{SIM}(3)$ pose is given, the problem mimics the mesh deformation problem introduced in Section III-B except that (7) is only the one-sided Chamfer distance and the optimization variable is the shape code instead of the vertices. These two observations lead to an intuitive alternating minimization algorithm to solve (7):
@@
After the pose step, the incremental $R$, $t$, and $s$ are accumulated onto the estimates $\hat{R}$, $\hat{t}$, and $\hat{s}$. Every time $\mathbf{c}$ is updated, the model points ($p_{n}$’s) are re-sampled to ensure sample uniformity as the areas of the mesh faces can change.
Shape step. In the shape step, we include the regularization losses (2) – (4). The final objective function $\mathcal{L}_{sc}$ (scsc: shape code) in the shape step is $\min_{\mathbf{c}}{\mathcal{L}_{sc}}=\min_{\mathbf{c}}\mathcal{L}_{ps}+\lambda_{n}\mathcal{L}_{n}+\lambda_{e}\mathcal{L}_{e}+\lambda_{l}\mathcal{L}_{l}$, where $\mathcal{L}_{ps}$ (psps: pose shape) refers to the objective function in (7) and the $\lambda$’s are the weights. These weights can be reused from Section III-B if we center and normalize the observed point cloud, and shift and normalize the transformed model points by the same amount (because the ShapeNet [35] objects in Section III-B are all centered and normalized). Since this is an iterative coordinate-descent style algorithm, we prefer small increments on $\mathbf{c}$. Only one step of gradient descent is performed on $\mathbf{c}$ during each shape step. However, we repeat the shape step several times in each iteration. The shape step is skipped in the last few iterations for better empirical performance.